# Computational Obfuscations and Random Oracles for Derandomizing Asynchronous Consensus

### v1

@@ 16       $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
This makes the task of the adversary simple: given $f$, it must find consecutive positions $v$ and $v+1$ such that $f(v)≠f(v+1)$, then steer the execution of the approximate agreement protocol to these values to prevent agreement. We aim to design $f$ to make this task computationally expensive relative to the costs paid by the processes. We assume that the primary cost to both the adversary and each process will be evaluating $f$ for particular inputs. Since each process evaluates $f$ exactly once, the main question becomes how many times the adversary evaluates $f$.
The worst case is when $f$ is a threshold function with an unknown threshold. Here, the adversary’s best strategy to find a bad pair $f(v)≠f(v+1)$ is a binary search, which requires $Θ(s)$ evaluations of $f$.
@@
Let $T$ be a threshold drawn uniformly from $\left\{{1,\dots,2^{s}}\right\}$, and let Let $f(v)=1$ if and only if $v≥T$. Then an adversary with oracle access to $f$ requires $Ω(s)$ evaluations of $f$ on average to find $v$ such that $f(v)≠f(v+1)$, while an adversary that is restricted to $s/2$ evaluations of $f$ finds such a $v$ with probability at most $2^{-s/2}$.
@@
Because $f$ is chosen without regard to the adversary’s strategy, we can limit our consideration to deterministic adversaries and then apply Yao’s Principle [20] to generalize to probabilistic adversaries. The rest of the argument is essentially the usual lower bound for binary search.
Fix some deterministic adversary strategy. We can express this strategy as a decision tree where each node specifies some value $v$ at which to compute $f(v)$, with two children corresponding to the result of observing $f(v)=0$ or $f(v)=1$. Each leaf of the tree is labeled by an output value $v$.
Observe that the adversary’s strategy succeeds in finding $v$ such that $f(v)≠f(v+1)$ precisely when $v=T-1$. So to succeed in all executions, the decision tree must have at least $2^{s}$ leaves to cover all possible values of $T$, which implies an average depth of at least $s$ by Kraft’s inequality. This gives the $Ω(s)$ lower bound to find a bad pair always.
If the tree has depth $s/2$, then it has at most $2^{s/2}$ leaves. The chance that a random $T$ lands on one of the at most $2^{s/2}$ values corresponding to one of these leaves is at most $2^{-s/2}$. ∎
The lemma immediately gives bounds on the cost of computing a bad schedule for Algorithm 5.1:
Corollary 5.3.
Assume that $f$ satisfies the conditions of Lemma 5.2, and let $t$ be a lower bound on the expected time complexity to compute $f(v)$. Then it costs $Ω(st)$ on average to find an execution of Algorithm 5.1 that violates agreement, and there is a constant $c>0$ such that any adversary limited to $cst$ total time finds an execution that violates agreement with probability at most $2^{-s/2}$.
In the following section, we show how to construct a function $f$ that satisfies the conditions of Lemma 5.2 and Corollary 5.3.
6 Computational obfuscation for threshold encoding
Specifically, we introduce a new approach for computationally obfuscating programs that use one-way function primitives. The program’s creation uses randomization that can be revealed by (tunable inversion processing required to) invert the one-way function. We further demonstrate the usefulness of such an obfuscation by enabling the execution of the obfuscated program, somewhat analogously to how programs operate on fully homomorphic encrypted data.
We start with the case in which the inversion of the one-way function can be easily performed, namely encrypting an array of values that is used to encode a threshold value $T$ (an array index) such that all entries that are less than $T$ encode 00 while the others encode $1$. Any asymmetric or symmetric encryption scheme (e.g., RSA, ElGamal, AES) can encrypt random strings that encode 00, say have bit parity 00, to be the first $T-1$ entries of the threshold encoding array, and the other entries to encode $1$. Then, provide the threshold encoding array and the decryption key as part of the computationally obfuscated program. The benefit of such obfuscation is more apparent when the computation needed to compute the obfuscated program is significantly smaller than the computation the adversary (in our case, the scheduler) needs to invest. Thus, we suggest using one-way functions that are employed in the context of proof of work, namely, SHA.
The $T-1$ first entries of the threshold encoding array are based on a random string $\langle choice,nonce,r,ed\rangle$, where choice is 00, while the the remaining $\ell-T+1$ entries have choice $1$. The value of $nonce$ is a random string chosen independently at random for each entry and appears in plaintext as an entry of a $nonce$ array in the program; the number of bits in the $nonce$ tunes the inversion difficulty. Using $nonce$ ensures that one (possibly exhaustive) search for an inverse is not helpful for an entry with a different $nonce$. $r$ is a random string, and $ed$ is an error detection code over the concatenated $nonce$ and $r$. $ed$ is used to avoid unplanned collisions of SHA that may lead to the existence of more than one $T$ value. A long enough $nonce$ can suffice to fulfill the $ed$ mission and allow to eliminate the $ed$ and use only $\langle choice,nonce,r\rangle$. In addition, $choice$ can be defined by the xor of all bits of $nonce$ and $r$, choosing a $nonce$ to fit the required 00 values in the $T-1$ prefix entries and $1$ values in the rest. In such a case, each entry of the threshold encoding array results from SHA applied to the concatenation of only the $nonce$ and $r$ associated with the entry.
Formally, let $H_{m}:\{0,1\}^{m}\rightarrow\{0,1\}^{m}$ be a secure hash function for some fixed $m\in\mathbb{N}$. In particular, we discuss the NIST post-quantum standard hash [18] SHA 512/256. Let $0<T≤\ell$ be a threshold value and let $f_{T}:\mathbb{Z}_{\ell}\rightarrow\{0,1\}$ be our threshold function for $T$, that is
$f_{T}(i)=\begin{cases}0&i<T\\ 1&i≥T\end{cases}$
1 procedure $\textnormal{{preprocess}}(\ell,T,m,k$ s.t. $t\leq\ell\wedge k<m)$

### v2

@@ 16 $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
If $f$ requires $T(f)$ time to evaluate, the time complexity incurred by each process running Algorithm 5.1 is $O(s)+T$. From Theorem 4.1 this implies that there is an adversary strategy that computes a bad execution in time $O(s^{2}+sT)$; this also tells us that at least one bad execution exists. We’d like to show a comparable lower bound on the cost to the adversary, assuming it is difficult to find a threshold in $f$.
Since every process has a value in $\left\{{v_{s},v_{s}+1}\right\}$, agreement holds automatically if $f(v_{s})=f(v_{s}+1)$. So an adversary that finds an execution that produces disagreement also finds a value $v_{s}$ with $f(v_{s})≠f(v_{s}+1)$.
For any fixed $f$ with $f(0)=0$ and $f(2^{s})=1$, there is a random-access machine that outputs a threshold value $v$ with f(v)f(v+1)f(v)≠f(v+1) in $O(1)$ time, since it can just include $v$ in its code. We will need to avoid this by considering the average-case cost to find a threshold value for an $f$ drawn from some family of threshold functions. Applying Algorithm 5.1 to the functions in this family then give a family of consensus protocols.
Theorem 5.2.
Let $F$ be a family of threshold functions with range $\left\{{0,\dots,\ell}\right\}$. Fix some some $n$ and some $s$ such that $2^{s}$$\ell$, and for each fFf∈F, let $P_{f}$ be the obfuscated consensus protocol given by instantiating the $s$-round version of Algorithm 5.1 with $f$ as its decision function.
Suppose we sample $P_{f}$ uniformly at random and consider a starting configuration in which not all processes have the same input. If there is a probabilistic random-access machine $M$ that takes $P_{f}$ and the process’s initial states as input, runs for $T$ expected time, and outputs a schedule that produces disagreement, then there is a probabilistic random access machine $M^{\prime}$ that takes $f$ as input, runs for $T+O(ns)$ expected time, and outputs a $v$ with f(v)f(v)+1f(v)≠f(v)+1.
Proof.
We give a construction for $M^{\prime}$.
Run $M$ on $P_{f}$ to obtain a schedule for Algorithm 5.1, using $T$ time. Simulate the algorithm until each process $p$ returns a value $i_{p}$ from approximateAgreement, but do not have any process call $f$. this takes $O(ns)$ time since each simulated step can be done in constant time in the RAM model. Output the smallest $i_{p}$ value as $v$.
The cost of this procedure is bounded by $T+O(ns)$. From Lemma 5.1 we know that all processes obtain either $v$ or $v+1$, and in the case that $M$ outputs a schedule that produces disagreement, we have f(v)f(v+1)f(v)≠f(v+1). ∎
In the following section, we give an implementation of a family of threshold functions with these properties.
6 Implementing an obfuscated threshold function
6.1 Definition
We introduce a new approach for computationally obfuscating programs that use one-way function primitives. The program’s creation uses randomisation that can be revealed by (tunable) inversion processing required to invert the one-way function. We construct the obfuscation by using a random oracle to hide a sequence of bits that encode the truth table of a threshold function. To obtain the output from the obfuscation, one must find the pre-image of a single hash. Thus it is imperative that the hardness of computing the pre-image be tunable such that it is feasible to find a single pre-image, yet infeasible for the computationally bounded adversary to enough pre-images to de-obfuscate the program. We achieve this tunability not by weakening the hash function, but by providing a prefix for the pre-image, such that the processes need only complete it. We begin with the following definitions.
Definition 6.1 (Threshold Function).
Let $T,\ell\in\mathbb{N}$ be two integers such that $0<T<\ell$. Define function $f_{T}:[\ell]\rightarrow\{0,1\}$ as
$f_{T}(i)=\begin{cases}0&i<T\\ 1&i\geq T\end{cases}$
Then $f_{T}$ is a threshold function for threshold $T$.
Definition 6.2 (Random Oracle Hash Function).
Let $m\in\mathbb{N}$ and let $H_{m}:\{0,1\}^{m}\rightarrow\{0,1\}^{m}$ be sampled uniformly at random from the set of all functions $\{0,1\}^{m}\to\{0,1\}^{m}$ (i.e., $H_{m}$ is a random oracle / random function). Then for all distinct $x\neq x^{\prime}\in\{0,1\}^{m}$ and all $y,y^{\prime}\in\{0,1\}^{m}$,
  1. 1.
    $\Pr\left[H_{m}(x)=y\right]=2^{-m}$
  2. 2.
    $\Pr\left[H_{m}(x)=y\ \wedge\ H_{m}(x^{\prime})=y^{\prime}\right]=2^{-2m}$
Definition˜6.2 is adapted from the book on cryptography by Katz and Lindell (KatzLindell ), Chapter 5.1. Using brute-force search, one must try at most $2^{m}$ different inputs to find a pre-image of a given output. The following corollary of Definition˜6.2 stems from this fact and states that given a prefix of the pre-image, our search space decreases.
Corollary 6.3 (Tunable pre-image runtime).
Let $H_{m}:\{0,1\}^{m}\to\{0,1\}^{m}$ be a random oracle (Definition˜6.2). Fix $k<m$ and a prefix $p\in\{0,1\}^{k}$, and let $H_{m}(p\|r)=y$ be the output of $H_{m}$ for some extension $r$ of $P$. For any (possibly randomized) algorithm $\mathcal{X}$ that makes at most $q$ oracle queries to $H_{m}$ and outputs $z\in\{0,1\}^{m-k}$, it holds that
$\Pr_{H_{m},\mathcal{X}}[H_{m}(p\|z)=y]\leq\frac{q}{2^{m-k}}$
Definition 6.4 (Threshold obfuscation scheme).
Let $\Pi=(Prep,Probe)$ be a pair of algorithms. If given a representation of a threshold function $f_{T}$ (6.1), $Prep$ produces output $O_{T}=Prep(f_{T})$ such that, for any given input $x$ of $f_{T}$, $Probe(O_{T},x)=f_{T}(x)$ holds with high probability, then $\Pi$ is a threshold obfuscation scheme.
We move one to formulate our security definition. We base this definition of the adversarial experiment format used in the book by Katz &\& Lindell KatzLindell . There exists a definition for obfuscation in DBLP:journals/jacm/BarakGIRSVY12 (Boaz Barak et al.). This definition requires that no probabilistic polynomial-time adversary can generate the same outputs as the original program with high probability. While we would like to have such obfuscation for the threshold function (a topic for possible future work), our current obfuscation construction does not satisfy this very strong notion; rather, we require that to reproduce the same outputs, an adversary must have a runtime that is at least logarithmic in the size of the construction, and an adversary must reproduce the exact outputs with probability $1$; otherwise, it cannot do any better than a random guess. Specifically, since we tailor this obfuscation to a threshold function, we formulate this specific notion of obfuscation as an experiment where the adversary gets the obfuscated threshold function just like the processes, and must recover the threshold $T$ in order to succeed.
Definition 6.5 (Obfuscation Security Experiment $Obf_{\mathcal{A},\Pi,\ell}$).
Let $\ell\in\mathbb{N}$ be the range of values for threshold $T$. Let $\Pi=(Prep,Probe)$ be an obfuscation scheme (6.4). Let $\mathcal{T}_{Probe}$ be the runtime of algorithm $Probe$. Let $\mathcal{A}$ be an adversarial algorithm with runtime $\mathcal{T_{A}}<\mathcal{T}_{Probe}\cdot\log\ell$.
  1. 1.
    A random $T\in[\ell]$ is chosen.
  2. 2.
    Compute $O_{T}=Prep(f_{T})$.
  3. 3.
    $\mathcal{A}$ is given $O_{T}$.
  4. 4.
    $\mathcal{A}$ returns an integer $T^{\prime}$.
  5. 5.
    Let $\tau$ be the number of times $\mathcal{A}$ computes $H_{m}$.
  6. 6.
    The output of the experiment is $1,\tau$, if $T^{\prime}=T$, otherwise the output is $0,\tau$.
Using Definition˜6.5, we formulate a definition of security for a threshold obfuscation scheme. While Katz&Lindell would define the security as an upper bound on the probability of success in the experiment $Obf$, in our case, the adversary must be able to cause the protocol to terminate with certainty. We therefore require that the number of hashes the adversary must compute in order to find $T$ with certainty is bounded in expectation.
Definition 6.6.
A threshold obfuscation scheme (Definition˜6.4) $\Pi=(Prep,Probe)$ is secure if for every adversary algorithm $\mathcal{A}$, it holds that
$\mathbb{E}\left[\tau\mid b=1\right]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
where $b,\tau=Obf_{\mathcal{A},\Pi,\ell}$.
While a bound on the probability of success would be stronger, this definition is sufficient to show that the adversary must do more work than the processes in order to cause the protocol to fail. We move on to define our threshold obfuscation scheme. We begin with the $Prep$ algorithm, the output of which will be used to compute the output of the threshold function in the $Probe$ algorithm. We call this algorithm preprocess.
preprocess 1. Fix integers $1<k<m$ and $\ell\geq 2$. 2. Choose threshold $1<T<\ell$ uniformly at random. 3. For all $1\leq i\leq\ell$, sample binary strings $P_{i}\leftarrow\{0,1\}^{k}$ and $r_{i}\leftarrow\{0,1\}^{m-k}$ uniformly at random. 4. For all $1\leq i\leq\ell$ (a) compute $d=H(P_{i}\|r_{i})$ (b) set $C_{i}\leftarrow d_{1}d_{2}\dots d_{m-1}$ that is, $C_{i}$ is the first m1m-1 bits of $D$ (c) set $V_{i}\leftarrow d_{m}\oplus f_{T}(i)$ 5. Return $(P_{1},\dots,P_{\ell},C_{1},\dots,C_{\ell},v_{1},\dots,v_{\ell})$.
Next, we define the $Probe$ algorithm, which we will use to compute a single output of $f_{T}$.
probe 1. Assume access to $H_{m}$ as well as $C_{1},\dots,C_{\ell}$, $P_{1},\dots,P_{\ell}$, and $v_{1},\dots,v_{\ell}$. 2. Let $1\leq i\leq\ell$ be an input of $f_{T}$. 3. Find string $r\in\{0,1\}^{m-k}$ such that $H_{m}(P_{i}\|r)]_{j}=[C_{i}]_{j}$ for all $j=1,\dots,m-1$. 4. Return $H_{m}(P_{i}\|r)]_{m}\oplus v_{i}$
A more complete definition and implementation can be found in Section˜6.3.
6.2 Correctness &\& security
We prove that Algorithms 6.1 and 6.2 (labeled as preprocess and probe) constitute a threshold obfuscation scheme (6.4), and also satisfy the security definition 6.6. In ˜6.8 we show the correctness of the construction, and in ˜6.11 we show that it is secure. We begin with the following lemma.
@@ 6.2 Correctness &\& security
[Distinct pre-image with prefix] Let $k,m\in\mathbb{N}$ be two integers where $k<m$, and let $H_{m}$ be a random oracle 6.2. Let $p\in\{0,1\}^{k},\quad r\in\{0,1\}^{m-k}$ be two binary strings. It holds that
$\Pr\left[\exists_{x\neq r}(H_{m}(p||x)=H_{m}(p||r))\right]\leq 2^{-k}$
@@
Let $p\in\{0,1\}^{k},\quad r\in\{0,1\}^{m-k}$ be two binary strings. Choose uniformly at random $x\in\{0,1\}^{m-k}\setminus\{r\}$. By Definition˜6.2, it holds that
$\Pr[H_{m}(p\|r)=H_{m}(p\|x)]\leq 2^{-m}$
By the union bound
$\Pr[\exists x(H_{m}(p\|r)=H_{m}(p\|x))]=\Pr[\bigcup_{x\neq r}H_{m}(p\|r)=H_{m}(p\|x)]$
$\leq\sum_{x\neq r}\Pr[H_{m}(p\|r)=H_{m}(p\|x)]$
Since there are $2^{m-k}-1$ possible choices for $x$ and assuming each sample of $x$ happens independently of the others,
$(2^{m-k}-1)\cdot\Pr[H_{m}(p\|r)=H_{m}(p\|x)]=(2^{m-k}-1)\cdot 2^{-m}\leq 2^{-k}$
Hence, we obtain that
$\Pr[\exists x(H_{m}(p\|r)=H_{m}(p\|x))]\leq 2^{-k}$
The following is the construction correctness claim. We require that the parameters $k,\ell$ satisfy $\frac{\ell}{2^{k}}=negl(k)$, meaning that their ratio is negligible in $k$. This assumption is reasonable since increasing $k$ decreases the probability of a collision as well as reduces the amount of work required to find the suffix of a pre-image.
Claim 6.8.
The algorithm pair $(\textnormal{{preprocess}},\textnormal{{probe}})$ is a threshold obfuscation scheme 6.4, for $\ell,k$ that satisfy $\frac{\ell}{2^{k}}=negl(k)$.
Proof.
Let $\ell\in\mathbb{N}$ be some positive integer and let $f_{T}$ be some threshold function where $0<T<\ell$. Fix some $m,k$ such that $0<k<m$. Let $C,P=\textnormal{{preprocess}}(\ell,T,m,k)$ be the outputs of algorithm preprocess. For each possible input $i\in[\ell]$ of $f_{T}$, the algorithm executes the following steps:
In lines 6.1, 6.1, the algorithm picks random strings $P[i]\in\{0,1\}^{k}$ and $r_{i}\in\{0,1\}^{m-k}$. Their concatenation $P[i]\|r_{i}$ is a string of length $m$ which is a valid input for $H_{m}$.
In 6.1, the algorithm computes the hash of $d\leftarrow H_{m}(P[i]\|r_{i})$.
In 6.1, the algorithm stores the first m1m-1 bits of $d$ in $C[i]$.
In 6.1, the algorithm stores $f_{T}(i)$ in $v_{i}$ by assigning $v_{i}\leftarrow d_{m}\oplus f_{T}(i)$.
Therefore, for any input $i$ of $f_{T}$, there exists at least one string $r_{i}\in\{0,1\}^{m-k}$ such that $H_{m}(P[i]\|r_{i})[j]=C[i][j]$ for all $j=1,\dots,m-1$ and $H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)=V[i]$.
Algorithm probe has access to $\ell,m,k,C,P,V$ and is given input $i$ of the threshold function. The algorithm finds suffix $y$ such that $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i][1,\dots,m-1]$. It remains to show that $H_{m}(P[i]\|y)[m]\oplus V[i]=f_{T}(i)$ holds with high probability. Assume the contrary.
This implies $H_{m}(P[i]\|y)[m]\oplus H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)\neq f_{T}(i)$ which can only happen if $H_{m}(P[i]\|y)[m]\neq H_{m}(P[i]\|r_{i})[m]$. Since $H_{m}$ is deterministic, this implies that $y\neq r_{i}$. But by Lemma˜6.7, the probability that this occurs is at most $2^{-k}$. Thus, for a large enough $k$, it holds that $y=r_{i}$ with high probability. Thus, $H_{m}(P[i]\|y)=H_{m}(P[i]\|r_{i})[m]$, which means that, w.h.p
$H_{m}(P[i]\|y)\oplus V[i]=H_{m}(P[i]\|r_{i})[m]\oplus H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)=f_{T}(i)$
Lemma 6.9.
For any adversary $\mathcal{A}$ that recovers $T$ from $C,P,V$, it holds that
$\mathbb{E}[\tau]\geq\frac{1}{2}\left((2^{m-k}+1)\cdot\log(\ell+1)\right)$
where $\tau$ is the number of times $\mathcal{A}$ computes $H_{m}$.
Proof.
By Definition˜6.2, given $C,P,V$, the threshold $T$ that is hidden in $C,P,V$ can be any value in $\{1,\dots,\ell\}$ with equal probability. Moreover, Definition˜6.2 implies that for each $i\in[\ell]$, every suffix $y\in\{0,1\}^{m-k}$ has equal probability of satisfying $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i]$. It follows that any adversarial algorithm that attempts to compute $T$ directly from $C,P,V$ must check $H_{m}(P[i]\|y)[1,\dots,m-1]=C[i]$ for some number of strings $y$. Order the set $\{0,1\}^{m-k}$ by some arbitrary order (by the number each string represents for example). Denote $X_{i,j}=H_{m}(P[i]\|y)$ where $y$ is the $j^{th}$ string in $\{0,1\}^{m-k}$. Denote $r=2^{m-k}$. Let $M\in\{0,1,\bot,?\}^{\ell\times r}$ be the matrix such that
[…]
Notice that:
  • (a)
    Since the adversary only needs to find one suffix $y$ per column $i$ that satisfies $X_{i,j}[1,\dots,m-1]=C[i]$, it holds that $M_{i,j}=1\vee 0\implies\forall_{g\neq j}M_{i,g}=\bot$.
  • (b)
    Since $C,P,V$ encode the truth table of a threshold function, the adversary can know that if $M_{i,j}=0$ then for all $1\leq r\leq i$ it holds that $M_{r,j^{\prime}}=0$ for some $j^{\prime}$, meaning $f_{T}(r)=0$.
We call the process of checking if a single suffix $y$ satisfies $X_{i,j}[1,\dots,m-1]=C[i]$ a probe of $M$ (not to be confused with the probeprobe algorithm). Denote $M^{(t)}$ as the state of $M$ after the adversarial algorithm makes $t$ probes. Conversely, denote $M^{(0)}$ as the initial state of $M$ at the start of the execution, this implies $M^{(0)}_{i,j}$=? for all $i,j$. Let $L_{t}$ be the number of columns in $M^{(t)}$ with at least one ?? cell (conversely, without any $1$ or 0 cells). Let $P_{t}$ be the number of non-?? cells, in columns with at least one ?? in $M^{(t)}$. Define potential function $\Phi$ as follows
$\Phi(L,P)=\frac{1}{2}\left((r+1)\cdot\log(L+1)-P\right)$
We show that $\Phi$ is the lower bound on the expected number of probes $\mathcal{A}$ must execute to find $T$.
Claim 6.10.
Denote $\text{hit}_{t}$ as the event where the $t^{th}$ probe $\mathcal{A}$ makes is at index $i,j$ and it holds that $X_{i,j}[1,\dots,m-1]=C[i]\wedge(X_{i,j}[m]\oplus V[i]=1\vee X_{i,j}[m]\oplus V[i]=0)$. Conditioned on $\text{hit}_{t}$, it holds that $\log(L_{t+1}+1)\geq\log(L_{t}+1)-1$, that is, a successful probe reduces the number ofpossibilities for $T$ by at most half.
Proof. Since $\mathcal{A}$ knows that $x=\max\{i\mid f_{T}(i)=0\}\leq T\leq\min\{i\mid f_{T}(i)=1\}$, it follows that $T$ is uniform on $\{x,x+1,\dots,x+L_{t}\}$, it follows that
$\Pr[f_{T}(x+j)=0]=\Pr[T>x+j]=\frac{j}{L_{t}+1},\qquad\Pr[f_{T}(x+j)=1]=\Pr[T\leq x+j]=\frac{L_{t}+1-j}{L_{t}+1}$
If $f_{T}(x+j)=0$ then $T\in\{x,\dots,x+j-1\}$ and the new interval size is $j$.
If $f_{T}(x+j)=1$ then $T\in\{x+j,\dots,x+L_{t}\}$ and the new interval size is $L_{t}+1-j$.
This implies that
$\mathbb{E}\left[\log(L_{t+1}+1)\mid\text{hit}_{t}\right]=\frac{j}{L_{t}+1}\log j+\frac{L_{t}+1-j}{L_{t}+1}\log(L_{t}+1-j)$
this expression is minizied at $j=\frac{L_{t}+1}{2}$, thus
$\mathbb{E}\left[\log(L_{t+1}+1)\mid\text{hit}_{t}\right]\geq\log((L_{t}+1)/2)=\log(L_{t}+1)-1$
$\mathbb{E}\left[\log(L_{t+1}+1)-\log(L_{t}+1)\mid\text{hit}_{t}\right]\geq-1\Diamond$
[…]
$\Pr[\text{hit}_{t}]=\frac{1}{r-q_{t}}$
If $\text{hit}_{t}$ does not occur, then $P_{t+1}=P_{t}+1$.
If $\text{hit}_{t}$ does occur, then by (b), we have that $P_{t+1}\leq P_{t}+1-(q_{t}+1)=P_{t}-q_{t}$.
Denote $\Delta P=P_{t+1}-P_{t}$, it follows that
$\mathbb{E}[\Delta P]\leq(1-\Pr[\text{hit}_{t}])+\Pr[\text{hit}_{t}]\cdot(-q_{t})=1-\Pr[\text{hit}_{t}]\cdot(q_{t}+1)$
$\mathbb{E}[\Delta P]\geq\Pr[\text{hit}_{t}]\cdot(q_{t}+1)-1\quad\text{(c)}$
Similarly, denote $\Delta\log(L+1)=\log(L_{t+1}+1)-\log(L_{t}+1)$ and $\Delta\Phi=\Phi(L_{t+1},P_{t+1})-\Phi(L_{t},P_{t})$. It follows that
$\mathbb{E}[\Delta\Phi]=\frac{1}{2}\left((+1)\mathbb{E}[\Delta\log(L+1)]-\mathbb{E}[\Delta P]\right)$
$\frac{1}{2}\left((+1)\Pr[\text{hit}_{t}]\cdot\mathbb{E}[\Delta\log(L+1)\mid\text{hit}_{t}]-\mathbb{E}[\Delta P]\right)$
By ˜6.10 and (c),
$\mathbb{E}[\Delta\Phi]\geq\frac{1}{2}\left((+1)\cdot\Pr[\text{hit}_{t}]\cdot(-1)-1+\Pr[\text{hit}_{t}]\cdot(q_{t}+1)\right)$
$\frac{1}{2}\cdot\left(-1+\Pr[\text{hit}_{t}](q_{t}-r)\right)$
$\frac{1}{2}\cdot\left(-1+\frac{1}{r-q_{t}}(q_{t}-r)\right)$
Thus
$\mathbb{E}[\Delta\Phi]\geq\frac{1}{2}(-1+(-1))=-1\quad\text{(d)}$
Having bounded the expected change in $\Phi$, we bound the expected number of steps $t$ that $\mathcal{A}$ will need to make in order to find $T$. We begin by showing that $\Phi+t$ is submartingale. Notice that when $L_{t}=0$ then the adversary knows all outputs of $f_{T}(\cdot)$, thus we define stopping time as
$\tau=\min\{t\mid L_{t}=0\}$
By definition, if $L_{\tau}=0$ then there are no unknown columns at time $\tau$; thus, $P_{\tau}=0$, which implies that
$\Phi(L_{\tau},P_{\tau})=\Phi(0,0)=\frac{1}{2}\left((+1)\log(0+1)-0\right)=0$
Let $\mathcal{F}_{t}=\sigma(M^{(t)})$ be the filtration generated by $M$ after $t$ probes. Denote $\Phi_{t}=\Phi(L_{t},P_{t})$. By (d), it holds that
$\mathbb{E}[\Phi_{t+1}-\Phi_{t}\mid\mathcal{F}_{t}]\geq-1$
$\implies\mathbb{E}[\Phi_{t+1}-\Phi_{t}+1\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+t)-(\Phi_{t}+t)+1\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+(t+1))-(\Phi_{t}+t)\mid\mathcal{F}_{t}]\geq 0$
Since $\Phi_{t}$ is completely determined by $M^{(t)}$, $\mathbb{E}[\Phi_{t}\mid\mathcal{F}_{t}]=\Phi_{t}$. This implies that
$\mathbb{E}[(\Phi_{t+1}+(t+1))-(\Phi_{t}+t)\mid\mathcal{F}_{t}]=\mathbb{E}[(\Phi_{t+1}+(t+1))\mid\mathcal{F}_{t}]-\mathbb{E}[(\Phi_{t}+t)\mid\mathcal{F}_{t}]\geq 0$
$\implies\mathbb{E}[(\Phi_{t+1}+(t+1))\mid\mathcal{F}_{t}]\geq\mathbb{E}[(\Phi_{t}+t)\mid\mathcal{F}_{t}]=\Phi_{t}+t\quad(e)$
Therefore, $\Phi_{t}+t$ is submartingale. Notice that for $\ell\cdot r$ probes, the adversary knows all columns of $M$ and the algorithm stops, thus it follows that $\tau\leq\ell\cdot r<\infty$ is bounded. Therefore, the optional stopping theorem applies to $\Phi_{t}+t$. By (e), and the optional stopping theorem
$\mathbb{E}[\Phi_{\tau}+\tau]\geq\Phi_{0}$
since $\Phi_{\tau}=0$
$\implies\mathbb{E}[\tau]\geq\Phi_{0}=\frac{1}{2}\left((2^{m-k}+1)\cdot\log(\ell+1)\right)$
The following is a security claim based on the established definition.
Claim 6.11.
The threshold function obfuscation $\Pi=\textnormal{{preprocess}},\textnormal{{probe}}$ is secure 6.6.
Proof.
Let $\mathcal{A}$ be any adversarial algorithm that tries to recover $T$ from $C,P,V$. Define $\tau,b=Obf_{\mathcal{A},\Pi,\ell}$ By Lemma˜6.9, if $\mathcal{A}$ finds $T$ and does $\tau$ probes to $H_{m}$, it holds that
$\mathbb{E}[\tau]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
therefore
$\mathbb{E}[\tau\mid b=1]\leq\frac{1}{2}\left((2^{m-k}+1)\log(\ell+1)\right)$
6.3 Implementation
Here we give a more detailed description of the obfuscated threshold implementation outlined in Section 6.1.
6.3.1 Outline
We present an implementation of the first algorithm of our threshold obfuscation scheme, Algorithm 6.1. This algorithm accepts hash size parameter $m$, hardness parameter $k$, value range $\ell$, and input threshold $T$. It returns three arrays, $C$, $P$ and $V$. The idea is to hide the threshold $T$ in an array of bits $V=v_{1},\dots,v_{\ell}$. We create an array of hashes $C=C_{1},\dots,C_{\ell}$ where each hash is created by randomly choosing two strings $p_{i}\in\{0,1\}^{k}$ and $r_{i}\in\{0,1\}^{m-k}$ for each $i=1,\dots,\ell$ and hashing their concatenation $p_{i}\|r_{i}$. We store in $C_{i}$ the first m1m-1 bits of the hash $H_{m}(p_{i}\|r_{i})$ and the last bit we XOR with the value of $f_{T}(i)$ to obscure it. We store this in the bit $v_{i}\leftarrow[H_{m}(p_{i}\|r_{i})]_{m}\oplus f_{T}(i)$. In this way, in order to compute $f_{T}(i)$, one must find a string $y$ such that $H_{m}(p_{i}\|y)]_{1\dots(m-1)}=C_{i}$, and then then compute $H_{m}(p_{i}\|y)]_{m}\oplus v_{i}$. Given that there is a high probability that $H_{m}(p_{i}\|y)]_{1,\dots,(m-1)}=[H_{m}(p_{i}\|r_{i})]_{1,\dots,(m-1)}\implies[H_{m}(p_{i}\|y)]_{m}=[H_{m}(p_{i}\|r_{i})]_{m}$ it follows that, $H_{m}(p_{i}\|y)]_{m}\oplus v_{i}=[H_{m}(p_{i}\|y)]_{m}\oplus[H_{m}(p_{i}\|r_{i})]_{m}\oplus f_{T}(i)=f_{T}(i)$ holds w.h.p. To create the hashes, we employ our random oracle function (6.2). Since one needs to find a suffix for a pre-image of the hash (up to the last bit) in order find $f_{T}(i)$, we must ensure that finding such suffix is feasible. To control the size of the search space we have a security parameter $k$ which determines the size of the prefix string $p_{i}$. For this reason, the preprocessing algorithm returns the array of prefix strings $P=P_{1},\dots,P_{\ell}$. Increasing $k$ decreases the size of the suffix required to find, which automatically reduces the search space. By the properties of a random oracle (6.2), reducing the size of the search space implies reducing the runtime required to find $r_{i}$.
In a practical implementation of the construction, the standard hash pub2015secure SHA 512/256 can be used in place of $H_{m}$ (Definition˜6.2). While no formal proof exists, it is widely accepted that in practice, SHA behaves like a random oracle, including an avalanche effect, where any minor change in the input, even a single bit, drastically alters the output.
1procedure $\textnormal{{preprocess}}(\ell,T,m,k$ s.t. $t\leq\ell\wedge k<m)$